The wavelength of light used in young's double slit experiment is 6000Å. The second and fourth bright fringes from the centre of the fringe pattern are located respectively at 10.24mm and 12.40mm, if the observation screen is placed at a distance of 1m from the slits. Calculate the separation b/w the slits.


1
Expert's answer
2021-06-14T17:27:16-0400

"x_2=\frac{2\lambda l}{d},"

"x_4=\frac{4\lambda l}{d},"

"d=\frac{n\lambda l}{x_n},"

"d=\frac{2\cdot 4\lambda l}{x_2\cdot x_4}=37.8~mm."